OpenStudy (anonymous):

A 0.89% (w/v) sodium chloride solution is referred to as physiological saline solution because it has the same concentration of salts as human blood. What is the molarity of a physiological saline solution.

OpenStudy (ookawaiioo):

1.5x10^-4 M

OpenStudy (anonymous):

Thanks.

OpenStudy (ookawaiioo):

remember 1000 g of water = 1 liter So 0.89% w/v = 8.9g of NaCl in 1000g of water or 1 liter of water

OpenStudy (anonymous):

Which I got. So taking 0.89mol/58.44g per mol equals 0.015mols. Right?

OpenStudy (ookawaiioo):

\[8.9\times \frac{ 1mol }{58g }= XXXX of moles \]

OpenStudy (ookawaiioo):

ooops ive made a mistake......

OpenStudy (ookawaiioo):

The answer is 1.5*10^-1 M ahh

OpenStudy (anonymous):

Yes! I got 0.15 M which is the same thing.

OpenStudy (ookawaiioo):

Sorry Im playing club music while i was doing this problem..... lol and GOOD JOB!

OpenStudy (anonymous):

Lol...awesome

OpenStudy (ookawaiioo):

You wanna become a doctor?

OpenStudy (anonymous):

yeah

OpenStudy (ookawaiioo):

how old are you?

OpenStudy (anonymous):

32

OpenStudy (ookawaiioo):

pm me Im in the same boat as you.

OpenStudy (anonymous):

awesome